Circular No. 6802

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

ECLIPSE OF LMC CLASSICAL CEPHEID D. Welch, Department of Physics and Astronomy, McMaster Univer-

sity, communicates: “The MACHO collaboration (cf. IAUC 6312) has discovered a 4.97-day-period classical cepheid in the Large Magellanic Cloud in an eclipsing-binary system the next primary minimum of which will occur on Feb. 25.0 (±3) UT. The system is located at α = 5^{h}20^{m}25^{s}.1, δ = −70^{o}11′08′′ (equinox 2000.0), has an orbital period of 400 days and an out-of-eclipse mean magnitude and color of V = 14.55, V − R = +0.07. The V semiamplitude of the system outside eclipse is ∼ 0.05 mag. The radial-velocity semiamplitude of the cepheid or its companion is likely ∼ 30 km/s, making the system useful for a determination of dynamical masses. Primary eclipse is between one and two weeks in total duration, during which the system becomes as faint as V = 15.2 and R = 15.1. It should be noted that the LMC will be progressively less well-placed for observing during primary eclipses in the following five years.” Potential observers are requested to contact Welch at welch@physics.mcmaster.ca, or by telephone at +1-905-525-9140 (extension 23186) to coordinate observations.

SUPERNOVA 1997ei IN NGC 3963 L. Wang, D. A. Howell and J. C. Wheeler, University of Texas, report:

“A spectrogram (range 420–813 nm; resolution ∼ 1 nm) of SN 1997ei was obtained with the 2.1-m telescope (+ IGP) at McDonald Observatory on 1997 Dec. 29. Preliminary analysis shows that the spectrum is similar to that of SN 1987M ∼ 1 week past maximum (Filippenko et al. 1990, A.J. 100, 1575) and SN 1994I close to maximum. This suggests that SN 1997ei may be of type Ic, rather than type Ia (IAUC 6796, 6800). The absorption features at 625 and 760 nm can be identified with Si II 635.5 nm and O I 777.4 nm, respectively. The O I 777.4-nm feature in SN 1997ei is stronger than the Si II 635.5-nm feature. The O I 777.4-nm feature in SN 1997ei is also stronger than the same feature in SNe 1994I and 1987M. Further observations are needed to confirm the classification. Our spectrum is accessible

at http://tycho.as.utexas.edu/~lifan/Images/sn97ei.gif.”

COMET 69P/TAYLOR CCD magnitude estimates: 1997 Oct. 8.79 UT, 17.8 (A. Sugie, Dynic

Observatory, 0.60-m reflector); 12.84, 18.3 (M. Yamanishi et al., Saji Observatory, 1.0-m reflector); Nov. 1.79, 17.9 (Sugie); 7.76, 17.6 (Yamanishi et al.); 9.78, 18.3 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector); Dec. 4.77, 16.4 (Nakamura); 24.71, 15.1 (Nakamura).

1998 January 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden